OpenStudy (anonymous):

Please help!!! What are the coordinates of the vertex of the function below? Write your answer in the form (x,y). Y+9=-6(x-2)^2

OpenStudy (kittiwitti1):

vertex form: y=a(x-h)^2+k vertex: (h,k)

OpenStudy (kittiwitti1):

isolate the y in that equation and then find the respective values for vertex coordinates (h,k)

OpenStudy (anonymous):

So what is the answer I'm lost!

OpenStudy (bee_see):

move the 9 to the right side. When you do that, the sign of 9 will change.

OpenStudy (kittiwitti1):

\[\text{subtracting y from both sides: }y+9-9=-6(x-2)^{2}-9\]\[\text{we now have }y=-6(x-2)^{2}-9\]\[\text{the vertex formula is }y=a(x-h)^{2}+k\]\[\text{Substitute the equation I solved for you into the vertex equation and find }(h,k)\]

OpenStudy (anonymous):

Ok I get it a little

OpenStudy (kittiwitti1):

Okay.

OpenStudy (kittiwitti1):

So as I said (sorry I deleted that), a = -6 We need to find h and k to get the vertex, (h,k) The formula is:\[y=a(x-h)^{2}+k\]We have a now, so\[y=-6(x-h)^{2}+k\]What else have you seen in the equation that matches the formula?

OpenStudy (kittiwitti1):

\[y=-6(x-2)^{2}-9\]\[y=-6(x-h)^{2}+k\]

OpenStudy (kittiwitti1):

What are the values for h and k, respectively?

OpenStudy (anonymous):

Is it 3 I'm really not good at math I'm trying

OpenStudy (bee_see):

look at the position of h and k...what number is over h? what number is over k?

OpenStudy (kittiwitti1):

It's okay, as long as you're trying :)

OpenStudy (anonymous):

-2,-9

OpenStudy (kittiwitti1):

Well... it's x-h, so x-(h).

OpenStudy (kittiwitti1):

x-h, x-2; so the answer is 2, not -2 (2,-9)

OpenStudy (anonymous):

I tried that answer that isn't it I'm taking a quiz online

OpenStudy (kittiwitti1):

Or you could look at it like this:\[y=a(x+(-h))^{2}+k\text{, finding the values of positive h and positive k}\] Hmm.

OpenStudy (bee_see):

You can look at this website for more examples: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

OpenStudy (kittiwitti1):

You mean 2,-9 is not the right answer? (We're actually not supposed to help you on quizzes, but just this once.)

OpenStudy (anonymous):

No it is not

OpenStudy (kittiwitti1):

Hmm. I just googled it and @satellite73 said it was the right answer in another question. Can you screenshot the quiz question for us?

OpenStudy (kittiwitti1):

@tamela12

OpenStudy (kittiwitti1):

Here is an example form if you need it: http://openstudy.com/study#/updates/5238b277e4b0b9d6b2d72ef8

OpenStudy (anonymous):

It want let me copy quiz I had same question and second time taking quiz

OpenStudy (kittiwitti1):

Here download this: https://app.prntscr.com/en/index.html

OpenStudy (bee_see):

does the problem have choices?

OpenStudy (kittiwitti1):

Maybe you put 2,-9 instead of (2,-9)?

OpenStudy (anonymous):

No choices

OpenStudy (kittiwitti1):

what I said xD

OpenStudy (anonymous):

Ok I'll see

OpenStudy (kittiwitti1):

>_>